OpenStudy (anonymous):

discuss the convergence of sequence where =(1/4+1) +(1/4^2+1) +(1/4^3+1) .....(1/4^n+1)

OpenStudy (anonymous):

@BSwan @ganeshie8

OpenStudy (anonymous):

I suppose n->infinity

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

it does not converge because of the plus one, essentially you have geometeric squenence of 1/4^n, which does converge, but the plus one to each term makes it inifinitly large, with an infininte number of plus ones, no convergence

OpenStudy (anonymous):

\(\large X_n = \sum_{n=1}^{n} \frac{1}{4^n} +1 \)

OpenStudy (anonymous):

so how do a theoreticaly prove it

OpenStudy (anonymous):

I*

OpenStudy (anonymous):

\(\large X_n = \sum_{n=1}^{n} \frac{1}{4^n} +\sum_{n=1}^{n} 1 \)

OpenStudy (anonymous):

\(\large X_n = \sum_{n=1}^{n} \frac{1}{4^n} + n \)

OpenStudy (anonymous):

\(\large \lim X_n =\lim \sum_{n=1}^{n} \frac{1}{4^n} +\ lim n \)

OpenStudy (anonymous):

\(\large \lim X_n = C +\lim n \)

OpenStudy (anonymous):

but lim n goes to + infinite

OpenStudy (anonymous):

so does not converge

OpenStudy (anonymous):

thanks ^^ again :P

OpenStudy (anonymous):

sry for laging :P so it not converge also not bounded hehe but its tend to infinity :D np , YW :)